A chemist wishes to mix a 10% salt solution with a 2% salt solution to obtain 6 liters of a 4% salt solution. How many liters of each should be added?
Answer by ptaylor(2052)


Let x=amount of 10% solution
And y=amount of 2% solution
Now we know that:
x+y=6------------------eq1
0.10x+0.02y=0.04*6--------------------eq2
substitute y=6-x from eq 1 into eq2 and we get
0.10x+0.02(6-x)=0.04*6 simplify
0.10x+0.12-0.02x=0.24 subtract 0.12 from both sides
0.10x+0.12-0.12-0.02x=0.24-0.12 collect like terms
0.08x=0.12 divide both sides by 0.08
x=1.5 l----------------------amount of 10% solution needed
y=6-x=6-1.5=4.5 l------------------amount of 2% solution needed
CK
1.5*0.10+4.5*0.02=0.04*6
0.15+0.09=0.24
0.24=0.24

Hope this helps---ptaylor